### v1

# O45 O1em m #3

Abstract
We introduce AG-VPReID, a challenging large-scale benchmark dataset for aerial-ground video-based person re-identification (ReID), comprising 6,632 identities, 32,321 tracklets, and 9.6 million frames captured from drones (15-120m altitude), CCTV, and wearable cameras. This dataset presents a real-world benchmark to investigate the robustness of Person ReID approaches against the unique challenges of cross-platform aerial-ground settings. To address these challenges, we propose AG-VPReID-Net, an end-to-end framework combining three complementary streams: (1) an Adapted Temporal-Spatial Stream addressing motion pattern inconsistencies and temporal feature learning, (2) a Normalized Appearance Stream using physics-informed techniques to tackle resolution and appearance changes, and (3) a Multi-Scale Attention Stream handling scale variations across drone altitudes. Our approach integrates complementary visual-semantic information from all streams to generate robust, viewpoint-invariant person representations. Extensive experiments demonstrate that AG-VPReID-Net outperforms state-of-the-art approaches on both our new dataset and other existing video-based ReID benchmarks, showcasing its effectiveness and generalizability. The relatively lower performance of all state-of-the-art approaches, including our proposed approach, on our new dataset highlights its challenging nature. The AG-VPReID dataset, code and models are available at https://github.com/agvpreid25/AG-VPReID-Net.
1 Introduction
Video-based person Re-Identification (ReID) is a challenging and in-demand task, with significant real-world applications in surveillance, search and rescue operations, and urban monitoring [25, 35, 37]. While traditional ReID methods focus on ground-based cameras [50, 7], the integration of aerial perspectives through aerial-ground person ReID presents a paradigm shift in this field [32]. This approach enables the identification and matching of individuals across non-overlapping aerial and ground-based camera views, substantially enhancing situational awareness and response times in complex environments [31, 46]. The motivation behind this research stems from the increasing deployment of aerial platforms, such as unmanned aerial vehicles (UAVs), which provide unique vantage points that complement ground-based observations. However, the development of robust aerial-ground ReID systems faces a significant challenge: the scarcity of diverse and large-scale datasets that capture the nuances of both aerial and ground perspectives. As demonstrated by ImageNet [4], large and diverse benchmarks are crucial for deep learning based methods, indicating a need for a comprehensive ReID dataset integrating multiple platforms, environments, and real-world challenges.
Dataset Year #Identities #Tracklets #Frames (M) #CV CC Att. Platform Dur. Altitude (m)
Ground Wearable Aerial
MARS [51] 2016 1,261 20,478 1.19 6 - -
LSVID [18] 2019 3,772 14,943 2.98 15 4 -
VCCR [10] 2022 392 4,384 0.15 1 90 -
CCVID [8] 2022 226 2,856 0.34 1 - -
MEVID [3] 2023 158 8,092 10.46 33 73 -
PDestre [17] 2020 253 1,894 0.10 1 - 5-6
G2A-VReID [47] 2024 2,788 5,576 0.18 2 - 20 - 60
AG-VPReID 2024 6,632 32,321 9.6 6 20 15 - 120
Table 1: Comparison of AG-VPReID with existing video-based person ReID datasets. Above: ground-based datasets, Below: aerial-based datasets. CV: Camera Views, CC: Clothes-Change, Att.: Attributes (Soft-biometrics annotations), Dur.: Duration (days).
Initial efforts in aerial-ground person ReID have focused primarily on image-based tasks. For instance, Nguyen et al. [31] pioneer this area by releasing the first aerial-ground ReID dataset, which includes images from one drone and one CCTV camera capturing 21,983 images of 388 identities. They later expand the dataset to 100,502 images of 1,615 individuals [32]. Recently, Zhang et al. [46] collect a synthetic dataset named CARGO, containing 108,563 images representing 5,000 identities, to complement real-world datasets. Within video-based tasks, Zhang et al. [47] collect a video-based dataset called G2A-VReID, which consists of 185,907 images and 5,576 tracklets from one drone and one CCTV camera, featuring 2,788 identities. Building on these advancements, G2A-VReID could be expanded compared to ground-based datasets like MARS [51], which includes 20,000 tracklets and 1.19 million frames from six cameras. While current aerial-ground datasets are valuable, increasing identity variation and environmental diversity would improve model robustness for real-world applications.
In light of this, we introduce AG-VPReID, a challenging large-scale benchmark dataset for Aerial-Ground Video-based Person ReID. AG-VPReID comprises 6,632 identities, 32,321 tracklets, and approximately 9.6 million frames, captured across multiple dates and times of day using a combination of three platforms: aerial drones operating at various altitudes (15-120m), stationary CCTV cameras and wearable mobile cameras. This dataset significantly surpasses existing video-based ReID benchmarks in terms of scale, diversity, and real-world applicability with the highest number of identities, the highest number of tracklets, the highest drone flying altitudes, and the most diverse platforms. The key characteristics of AG-VPReID include: drastic view changes between aerial and ground perspectives; a large number of annotated identities across multiple sessions; rich outdoor scenarios with varying environmental conditions; significant differences in resolution between aerial and ground footage; and both controlled scenarios with clothing changes and in-the-wild pedestrian traffic.
Aerial-ground person re-identification presents unique challenges due to significant appearance variations between aerial and ground-level views. These variations include extreme viewpoint differences, drastic changes in resolution and scale, partial occlusions, and temporal discontinuities caused by high-flying altitudes and long-range captures. Traditional video-based person ReID methods, although effective in ground-based settings [26, 45], often struggle in aerial-ground scenarios due to the complex combination of inconsistent motion patterns and the aforementioned variations.
To address these limitations, we introduce AG-VPReID-Net, an end-to-end framework for Aerial-Ground Video-based Person Re-Identification. Unlike existing state-of-the-art methods focused on single-view or ground scenarios, AG-VPReID-Net features three complementary streams tailored for aerial-ground challenges: i) An Adapted Temporal-Spatial Stream enhances traditional temporal modeling by integrating identity-specific memory and temporal shape analysis. This improves the extraction of consistent motion patterns and body shape representations, addressing the temporal discontinuity and motion inconsistencies of aerial footage, outperforming standard LSTM [6, 14] or 3D CNN [20] approaches; ii) A Normalized Appearance Stream addresses the resolution and appearance differences between aerial and ground views by using UV maps aggregation across frames for a normalized appearance representation. This provides robustness against pose changes, viewpoint shifts, and varying image quality, excelling where current appearance-based methods [29, 19] falter; and iii) A Multi-Scale Attention Stream addresses scale variations inherent in aerial-ground data by incorporating multi-scale feature extraction, motion analysis, temporal context, and a transformer decoder, effectively improving identification across drone altitudes compared to single-scale [13, 43] methods. By integrating these streams, AG-VPReID-Net offers incremental improvements in aerial-ground video-based re-identification, highlighting its potential in addressing this challenging scenario.
In summary, our main contributions are as follows:
(1) We introduce AG-VPReID, a challenging large-scale benchmark for aerial-ground video-based person ReID, bridging the gap with a diverse dataset that captures nuanced challenges from both aerial and ground perspectives.
(2) We propose AG-VPReID-Net, an innovative end-to-end framework that integrates adapted temporal-spatial processing, normalized appearance representation, and multi-scale attention mechanisms to effectively address the challenges of aerial-ground ReID.
(3) AG-VPReID-Net sets new state-of-the-art performance on the AG-VPReID and existing video-based ReID benchmarks, demonstrating our approach’s effectiveness and generalizability across different settings.
Figure 1: Our dataset was captured using six diverse cameras (aerial drones, CCTVs, GoPros) with sample images and camera locations shown on the right. Left shows cross-camera appearance variations of two pedestrians across different sessions and times of the day.
2 Prior Work
Video-based Person ReID Datasets. Existing person ReID datasets are numerous but severely lack the ability to address real-world challenges, particularly in aerial-ground scenarios. Ground-based datasets like MARS [51] and LS-VID [18] provide large-scale benchmarks but focus mainly on ground perspectives, highlighting the need for multi-platform surveillance datasets. The inclusion of clothing changes in datasets such as MEVID [3], CCVID [8], and VCCR [10] represents progress in addressing real-world challenges, though they could benefit from more identities and diverse environments. The BRIAR dataset [2], while featuring 1,000 subjects and UAV footage, primarily targets face recognition with restricted access. P-Destre [17] pioneered aerial view exploration, though its use of a single drone at lower altitudes (5-6m) creates opportunities for datasets covering higher operational altitudes more common in surveillance applications. The G2A-VReID dataset by Zhang et al. [47] represents an important step in combining aerial and ground views. While innovative, it contains 2,788 identities within a 20-60m altitude range using 2 cameras, suggesting opportunities for future datasets to expand in scale, altitude diversity, camera count, and environmental variety. Tab. 1 compares our AG-VPReID dataset with others across multiple dimensions.
Video-based Person ReID Approaches. Video-based person ReID methods have evolved to leverage both spatial and temporal cues. Early approaches used recurrent neural networks and 3D convolutional networks [30, 20], while later works incorporated temporal pooling [51] and attention mechanisms [25]. Recent advancements include temporal complementary learning [13], Transformer-based architectures [28, 11], and techniques addressing cross-platform and cloth-changing scenarios [47, 39, 10]. The AG-ReID 2023 Challenge [33] highlighted aerial-ground ReID challenges, with winners employing re-ranking, data augmentation, and centroid-based representations. Recent works include Instruct-ReID [12] with instruction-guided retrieval, Domain Shifting [15] for distribution adaptation, and SEAS [53] using 3D body shape guidance. Despite these developments, most existing methods employ uni-modal frameworks trained on predefined label sets. In contrast, recent work has proposed a visual-language multi-modal learning paradigm [45], potentially offering more flexibility and robustness in feature representation for video-based person ReID.
3 AG-VPReID Dataset
This section provides an overview of our AG-VPReID dataset creation process. We describe our methods for collecting video footage in Sec.  3.1. Sec.  3.2 outlines our annotation procedures. Sec. 3.3 compares AG-VPReID with existing datasets, emphasizing its unique features.
3.1 Dataset Collection
The AG-VPReID dataset was collected over 20 days, comprising 10 morning sessions (8:30am-10:00am) and 10 afternoon sessions (3:30pm-5:00pm), each lasting 60 minutes. We captured data using two drones, two CCTV cameras, and two wearable cameras. Drones flew at four altitudes (15m, 30m, 80m, 120m) for 15 minutes each per session. The dataset contains 240 hours of video footage in total. Tab. 2 details the equipment specifications. The dataset includes varying resolutions, frame rates, and perspectives, ranging from ground level to 120-meter aerial views. Fig. 1 shows example frames across different viewpoints and image qualities from different platforms. The drones, CCTV, and wearable devices are positioned to view individuals from different angles, as illustrated in Fig. 1, forcing person ReID models to learn robust multiview and partial-view representations to be effective.
Type Model Resolution Lens FPS
CCTV Bosch (Outdoor) $704\times 480$ 24mm 15
Bosch (Indoor) $1280\times 720$ 18mm 25
Wearable GoPro10 (Front) $3840\times 2160$ 16mm 30
GoPro10 (Side) $1920\times 1080$ 16mm 60
Drones DJI Inspire2 $3840\times 2160$ 24mm 25
DJI M300RTK $8192\times 5460$ 35mm 1
Table 2: Equipment specifications for the AG-VPReID dataset.
To ensure professional drone operations, we hired a specialized team consisting of one Remotely Piloted Aircraft Systems (RPAS) engineer, one Chief Remote Pilot, and one RPAS technician. This team was present for all 20 data collection days, responsible for setting up and flying the drones, capturing high-quality aerial footage, and performing initial data preprocessing. Their expertise was crucial in maintaining consistency and quality across the dataset, particularly for the challenging aerial perspectives.
3.2 Labeling Process
The AG-VPReID dataset uses YOLOv8x for person detection and tracking [16], extracting images from all frames across multiple cameras. It includes both short-term and long-term ReID scenarios, with participants changing clothing for long-term testing. Identity matching was performed by multiple expert annotators with research assistant help to ensure accuracy and consistency. Following [32], we manually labeled 15 selective soft-biometrics attributes per identity to enhance attribute-based person ReID capabilities. For the attribute list, see 8.1.
3.3 Dataset Characteristics
Compared with existing video-based ReID datasets, our AG-VPReID dataset has five unique characteristics:
1) The highest number of identities and tracklets. With 6,632 unique identities and 32,321 tracklets, our dataset significantly exceeds existing ones. It has almost double the identities of ground-based datasets like LSVID [18] (3,772) and over six times the tracklets of G2A-VReID [47] (5,576). With 9.6 million frames, it provides 50 times more data than other aerial-ground datasets, though MEVID [3] has slightly more frames but fewer identities and tracklets.
2) The most diverse platforms. Our AG-VPReID dataset is the first to incorporate aerial, ground, and wearable platforms for video-based person ReID. The inclusion of wearable cameras introduces a novel dimension with high-quality first-person perspectives. This combination results in extreme variations in resolution and subject size across platforms: UAV (18$\times$37 to 293$\times$542 pixels), CCTV (22$\times$23 to 172$\times$413 pixels), and GoPro (25$\times$48 to 483$\times$1078 pixels).
3) The highest flying altitudes. AG-VPReID features footage from altitudes up to 120 meters, exceeding previous datasets’ 60-meter maximum [47]. This introduces challenges: 1) extreme viewpoints with perspective distortions; 2) multiple scales with varying resolutions; and 3) image quality issues (Fig. 2). We used two drones—one with a wide-angle camera for area monitoring and another with a narrow-angle camera for detailed observation.
4) Rich outdoor scenarios with real-world challenges. Our AG-VPReID dataset presents diverse outdoor campus scenarios with real-world challenges including complex occlusions, varied poses from different activities, and uniform-wearing individuals. Fig. 2 shows examples of these diverse scenarios.
5) Other notable characteristics. AG-VPReID includes comprehensive attributes for each identity (gender, age, clothing style, accessories), enabling fine-grained analysis. The dataset features long-term tracking data of 14 diverse individuals recorded across multiple days, each wearing different clothing per session to capture real-world variations. We also provide camera calibration information and GPS coordinates to support multi-camera tracking research. See supplementary material Sec. 8 for details.
Figure 2: Main challenges in AG-VPReID: extreme viewpoints, varying resolutions and sizes, pose/illumination variations, occlusions, and similar clothing.
3.4 Ethics and Privacy
The research has been ethics-approved by related authorities for data collection and usage of the collected data for research purposes. We use “Deface” [5] to blur facial regions in media to protect individual identities. Data is securely stored with authorized access only. For volunteers involved in the study, ethical clearance was obtained, and informed consent was secured prior to their participation. Details on privacy protection and data usage are available at our project repository.
4 Approach
We propose AG-VPReID-Net, a purpose-built framework addressing aerial-ground ReID’s unique challenges. In particular, we propose an Adapted Temporal-Spatial Stream for robust temporal-spatial representations to deal with the temporal discontinuity challenge caused by drone motion from unstable tracking between frames. We propose a Normalized Appearance Stream for resolution and appearance changes to deal with extreme viewpoint shifts. To deal with altitude-driven scale variance, we introduce a Multi-Scale Attention Stream for scale variations. Fig. 3 illustrates our architecture. Detailed stream contributions are provided in Tab. 9 of the supplementary material.
Figure 3: The three-stream AG-VPReID-Net architecture addresses aerial-ground ReID challenges: Temporal-Spatial stream for motion modeling and temporal features, Normalized Appearance for resolution/appearance variations, and Multi-Scale Attention for aerial-ground scale variations.
4.1 Stream 1: Adapted Temporal-Spatial Stream
When performing video-based person ReID, a key challenge is handling inconsistent motion patterns and temporal gaps between video frames. To address this, we propose an Adapted Temporal-Spatial stream that combines CLIP’s visual encoder with temporal and 3D shape modeling to create a comprehensive representation of individuals. Our method operates on a sequence $\mathcal{V}$ of $T$ frames through the following components:
Visual Feature Extraction: Using CLIP’s visual encoder $E_{v}(\cdot)$, we extract frame-level features,
$F_{t}$ $=E_{v}(\mathcal{I}_{t}),\quad t\in\{1,...,T\},$ (1)
where $\mathcal{I}_{t}$ and $F_{t}$ are the $t$-th frame and its features.
Temporal Processing: We incorporate temporal modeling through two key components:
1) Temporal 3D Shape Modeling (TSM): Following  [34], we extract 3D shape representations,
$g_{t}=\text{GRU}(F_{t},g_{t-1}),\beta_{t}=\text{3D Regressor}(g_{t}),$ (2)
where $g_{t}$ captures temporal dynamics and $\beta_{t}$ represents SMPL model parameters.
2) Temporal Feature Enhancement (TFE): Adapting from [45], we enhance features by combining appearance and shape,
$F_{enhanced}$ $=\text{TFE}(F_{1:T},\beta_{1:T}).$ (3)
Identity-Aware Processing: We incorporate identity information through,
$\mathcal{M}_{y_{i}}$ $=\frac{1}{N_{y_{i}}}\sum_{V\in\mathcal{V}_{y_{i}}}\text{TAP}(F_{enhanced}),$
$\mathcal{M}_{y_{i}}^{refined}$ $=\text{SSP}(F_{enhanced},\mathcal{M}_{y_{i}}),$
$R_{stream1}$ $=[F_{enhanced};\mathcal{M}_{y_{i}}^{refined}],$ (4)
where $\mathcal{M}_{y_{i}}$ is the identity memory bank constructed through Temporal Average Pooling (TAP), and the Sequence-Specific Prompt (SSP) module refines this representation for the final output $R_{stream1}$.
4.2 Stream 2: Normalized Appearance Stream
The Adapted Temporal-Spatial (ATS) stream provides robust temporal-spatial representation, but may not fully capture fine-grained appearance details across viewpoints, especially in aerial footage. To address this limitation, we propose a Normalized Appearance (NA) stream that effectively aggregates appearance details from multiple viewpoints.
The NA stream normalizes and combines appearance information across frames using UVTexture maps and visibility masks. Our process involves: (1) Extracting UVTexture maps and visibility masks per frame, (2) Normalizing UVTexture maps brightness, (3) Aligning maps across frames, (4) Weighted aggregation using visibility masks, and (5) Generating the final normalized representation. The brightness normalization and weighted aggregation of UVTexture maps can be formulated as,
$T_{i}^{norm}$ $=\gamma(H(N(T_{i}))),$ (5)
$T_{aggregated}$ $=\frac{\sum_{i=1}^{N}V_{i}\odot T_{i}^{norm}}{\sum_{i=1}^{N}V_{i}},$ (6)
where $T_{i}^{norm}$ is the normalized UVTexture map for frame $i$, $N(\cdot)$, $H(\cdot)$, and $\gamma(\cdot)$ are normalization, histogram matching and gamma correction functions respectively. $T_{aggregated}$ is the final aggregated map. We leverage PhysPT [49] for pose estimation and Texformer [41] to generate UV maps from PhysPT’s output 3D meshes. The maps are improved through inter-frame consistency before feeding into the DGC Omni-scale Module [52].
4.3 Stream 3: Multi-Scale Attention Stream
While the ATS stream provides a robust temporal-spatial representation and the NA stream addresses viewpoint and occlusion challenges, aerial-ground person ReID still faces significant hurdles due to extreme scale variations between drone and ground-level footage. The first two streams effectively capture temporal dynamics, 3D shape information, and viewpoint-invariant appearance details, but they may not fully address the drastic scale differences inherent in aerial-ground scenarios. To complement the ATS stream and NA stream and address this limitation, we propose a Multi-Scale Attention (MSA) stream.
In detail, this stream leverages the power of frozen large vision models combined with lightweight, adaptive processing. Specifically, this stream utilizes a frozen large vision model to extract multi-scale features for video-based person ReID. By combining a lightweight Transformer decoder with a local temporal module, this approach dynamically integrates spatial and temporal information, thereby enhancing our framework’s ability to accurately capture essential person-specific details.
Specifically, for each frame $\mathcal{I}_{t}$ within the sequence $\mathcal{V}{y_{i}}$, the CLIP vision encoder CLIP [36] is employed to extract features independently. The process collects tokens from various layers at regular intervals to compile a detailed feature map that captures spatial correspondences. These frame feature maps are subsequently concatenated and assembled into a spatiotemporal feature volume $\mathbf{G}$. Following the methods [23, 44], we integrate temporal information into this volume before processing it through a Transformer decoder. This decoder globally aggregates features across multiple layers, employing a video-level classification token as a query, with feature volumes from different layers of the backbone serving as keys and values. A linear layer then maps the output of the decoder’s final block to produce class predictions. The operational dynamics of the Transformer decoder are outlined as follows,
$Y_{i}$ $=\operatorname{Temp}_{i}([\mathbf{G}_{N-M+i,1},\mathbf{G}_{N-M+i,2},...,\mathbf{G}_{N-M+i,T}]),$
$\tilde{q}_{i}$ $=q_{i-1}+\operatorname{MHA}_{i}(q_{i-1},Y_{i},Y_{i}),$
$q_{i}$ $=\tilde{q}_{i}+\operatorname{MLP}_{i}(\tilde{q}_{i}),$
$f_{G}$ $=\operatorname{FC}(q_{M}),$ (7)
where $\mathbf{G}_{n,t}$ represents the features of frame $t$ extracted from the $n$-th layer of CLIP vision encoder. The feature volume $Y_{i}$, which undergoes temporal modulation, is input into the $i$-th layer of the Transformer decoder. The query token $q_{i}$ is incrementally refined, beginning with $q_{0}$ as learnable initial parameters. The final output $f_{G}$, corresponds to the final feature. The spatiotemporal decoder comprises $M$ blocks. $N$ denotes the number of encoder layers. Multi-head attention (MHA) involves query, key, and value, each of which plays a distinct role. The operator $\operatorname{Temp}(\cdot)$ is utilized to model temporal dynamics, which produces feature tokens influenced by detailed temporal information.
5 Experimental Results
5.1 Datasets and Evaluation Metrics
We evaluate our method on AG-VPReID and four established video-based person ReID datasets: iLIDS [38], Mars [51], LS-VID [18] and G2A-VReID [47]. For AG-VPReID, 3,013 identities with ground and aerial views were split 1:1 for training and testing [31, 32]. Training and testing details are presented in Tab. 3. We use Cumulative Matching Characteristic (CMC) at Rank-1 and mean Average Precision (mAP) for evaluation.
Case Subset # IDs # Tracklets # Images (M)
Training All 1,555 13,300 3.85
All 1,456 13,566 3.94
15m 506 4,907 1.50
Testing (A2G) 30m 377 2,885 0.89
80m 356 2,592 0.69
120m 308 3,182 0.86
All 5,075* 19,021 5.79
15m 1,403 6,362 2.14
Testing (G2A) 30m 1,406 4,468 1.41
80m 1,162 3,866 1.13
120m 1,195 4,325 1.11
Table 3: Statistics of AG-VPReID dataset. A2G: aerial-to-ground, G2A: ground-to-aerial. *3,619 additional IDs as distractors.
MARS LS-VID iLIDS-VID G2A-VReID AG-VPReID
Method Ground $\rightarrow$ Ground Ground $\rightarrow$ Aerial Aerial $\rightarrow$ Ground Ground $\rightarrow$ Aerial
mAP Rank-1 mAP Rank-1 Rank-1 Rank-5 mAP Rank-1 mAP Rank-1 mAP Rank-1
STMP[29] 72.7 84.4 39.1 56.8 84.3 96.8 - - 50.7 60.3 45.2 55.8
M3D[20] 74.1 84.4 40.1 57.7 74.0 94.3 - - 52.4 62.6 47.9 57.3
GLTR[19] 78.5 87.0 44.3 63.1 86.0 98.0 - - 55.6 65.8 50.1 60.5
TCLNet[13] 85.1 89.8 70.3 81.5 86.6 - 65.4 54.7 57.2 67.9 52.7 62.4
MGH[43] 85.8 90.0 61.8 79.6 85.6 97.1 76.7 69.9 60.3 70.8 55.5 65.2
GRL[25] 84.8 91.0 - - 90.4 98.3 - - 58.7 68.4 53.9 63.6
BiCnet-TKS[14] 86.0 90.2 75.1 84.6 - - 63.4 51.7 59.8 69.2 54.3 64.7
CTL[24] 86.7 91.4 - - 89.7 97.0 - - 56.4 66.9 51.8 61.3
STMN[6] 84.5 90.5 69.2 82.1 - - 66.7 56.1 61.6 71.5 56.9 66.2
PSTA[39] 85.8 91.5 - - 91.5 98.1 - - 60.5 70.2 55.8 65.7
DIL[11] 87.0 90.8 - - 92.0 98.0 - - 61.2 70.9 56.3 66.1
STT[48] 86.3 88.7 78.0 87.5 87.5 95.0 - - 61.0 70.7 56.1 65.9
TMT[28] 85.8 91.2 - - 91.3 98.6 - - 60.8 70.5 55.9 65.8
CAVIT[40] 87.2 90.8 79.2 89.2 93.3 98.0 - - 61.4 71.1 56.5 66.3
SINet[1] 86.2 91.0 79.6 87.4 92.5 - - - 61.3 71.0 56.4 66.2
MFA[9] 85.0 90.4 78.9 88.2 93.3 98.7 - - 61.1 70.8 56.2 66.0
DCCT[27] 87.5 92.3 - - 91.7 98.6 - - 61.5 71.2 56.6 66.4
LSTRL[26] 86.8 91.6 82.4 89.8 92.2 98.6 - - 61.7 71.3 56.7 66.5
CLIP-ReID[21] 88.1 91.7 80.6 88.8 - - - - 62.3 71.6 57.2 66.8
AG-VPReID-Net 91.5 93.2 87.3 93.2 96.3 99.5 81.3 73.1 64.0 71.9 58.0 75.6
Table 4: Performance comparison across datasets. Bold shows best results.
5.2 Implementation Details
Our pipeline leverages UV maps generated by Texformer [41] using 3D human meshes from PhysPT [49] with refined pose estimation. The UV maps are processed through normalization, histogram matching, and gamma correction before weighted blending with visibility masks. The architecture consists of three streams: an Adapted Temporal-Spatial Stream (CLIP ViT-B/16), a Normalized Appearance Stream for 3D coordinates and UV textures, and a Multi-Scale Attention Stream (CLIP ViT-L/14). Implementation details are in the supplementary material Sec. 10.
5.3 Comparison with State-of-the-Art Methods
We evaluate our proposed method AG-VPReID-Net against several state-of-the-art approaches across multiple video-based person ReID datasets. Tab. 4 summarizes the results.
Ground-to-Ground Datasets. Our method achieves superior performance on MARS (91.5% mAP, 93.2% Rank-1), exceeding CLIP-ReID by 3.4% mAP. On LS-VID (87.3% mAP, 93.2% Rank-1), we surpass LSTRL by 4.9% mAP. For iLIDS-VID, we reach 96.3% Rank-1, outperforming MFA by 3.0%.
Cross-Platform Datasets. On G2A-VReID, we achieve 81.3% mAP and 73.1% Rank-1, surpassing MGH by 4.6% mAP. Note that the G2A-VReID dataset only provides a ground-to-aerial testing set. For AG-VPReID, we demonstrate strong results in both Ground-to-Aerial (58.0% mAP, 75.6% Rank-1, exceeding CLIP-ReID by 8.8% Rank-1) and aerial-to-ground scenarios (64.0% mAP, 71.9% Rank-1, surpassing CLIP-ReID by 1.7% mAP and 0.3% Rank-1).
Aerial $\rightarrow$ Ground Ground $\rightarrow$ Aerial
Method Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10
St-1 71.52 80.42 83.88 74.80 84.27 86.90
St-2 58.40 70.20 75.80 61.50 73.60 78.20
St-3 61.65 74.53 79.15 67.38 78.82 82.3
St-12 69.50 78.80 82.50 72.80 82.60 85.40
St-13 71.80 80.60 83.91 75.40 84.48 86.91
St-23 65.70 76.55 80.90 70.10 80.45 83.91
St-123 71.91 80.67 83.92 75.57 84.50 86.92
Table 5: Ranking accuracy (%) improvement on AG-VPReID dataset.
Aerial $\rightarrow$ Ground Ground $\rightarrow$ Aerial
Method 15m 30m 80m 120m 15m 30m 80m 120m
St-1 80.28 78.76 67.24 52.47 83.25 83.03 67.07 62.31
St-2 69.75 68.13 52.62 38.12 72.87 71.41 52.22 45.43
St-3 74.25 74.01 52.53 35.13 77.18 78.91 56.59 52.56
St-12 78.32 76.82 65.24 50.53 81.34 81.27 65.17 60.43
St-13 80.55 78.95 67.50 52.85 83.70 83.55 67.60 63.10
St-23 76.45 74.90 57.40 42.70 79.55 79.40 57.40 52.80
St-123 80.66 79.00 67.63 53.00 83.92 83.66 67.82 63.32
Table 6: Rank-1 accuracy (%) on AG-VPReID at various altitudes.
5.4 Ablation Study
We conduct an ablation study on AG-VPReID to evaluate each stream. St-1 is our temporal modeling stream, St-2 is the appearance normalization stream, and St-3 is the multi-scale feature stream. Their combinations (St-12/13/23/123) merge multiple streams.
Stream Contributions. Tab. 5 shows St-1 achieves the strongest individual performance (71.52% A2G, 74.80% G2A Rank-1). St-2 and St-3 show moderate results (58.40% and 61.65% A2G Rank-1). Combined streams demonstrate complementary strengths, with St-123 achieving the best results (71.91% A2G, 75.57% G2A Rank-1) by integrating all three streams.
Figure 4: Baseline vs our method on AG-VPReID dataset. Green/red: correct/incorrect labels. First tracklet image shown. Ranks show improvements in bold.
Impact of Altitude. Table 6 shows performance decreases with altitude, dropping most between 30m and 80m. A2G Rank-1 drops  11% across streams due to scale variations and reduced resolution. St-1 shows remarkable robustness at high altitudes versus other streams (52.47% vs 38.12%/35.13% at 120m A2G), demonstrating a +17.34% improvement through our temporal modeling approach. Similarly, St-2’s physics-informed UV mapping provides a +7.57% Rank-1 improvement at 120m (42.70% vs. 35.13%), while St-3’s multi-scale attention design yields a +17.72% improvement (52.85% vs. 35.13%) in the same setting. The poor high-altitude performance by all methods highlights unique challenges not present in typical low-altitude datasets. The St-123 maintains best performance by leveraging complementary stream strengths across all altitudes.
Clothing Changes vs. Camera Angles. Analysis shows altitude increases (15m → 120m) reduce Rank-1 by 27.66%, significantly more than clothing changes (7.85% in ground-to-ground). Without clothing changes, aerial-ground matching (71.91% Rank-1) still underperforms ground-to-ground (91.52%) due to viewpoint differences. When combining aerial views with clothing changes (65.83% Rank-1), these factors create synergistic challenges where viewpoint differences amplify clothing ambiguity. See Table 7.
Table 7: Impact of clothing changes (CC) vs. camera angles.
Scenario Rank-1 mAP Key Observation
Camera Angle Impact
15m altitude (AG) 80.66 77.23 Baseline performance
30m altitude (AG) 79.00 75.81 Minimal degradation
80m altitude (AG) 67.63 63.42 $-13.03\%$ Rank-1 vs. 15m
120m altitude (AG) 53.00 48.75 $-27.66\%$ Rank-1 vs. 15m
Clothing Change Impact
GG-SameClothes 91.52 88.74 Upper-bound performance
GG-DiffClothes 83.67 79.92 $-7.85\%$ Rank-1 (CC-only impact)
AG-SameClothes 71.91 64.00 $-19.61\%$ Rank-1 (AG-only impact)
AG-DiffClothes 65.83 57.52 $-6.08\%$ Rank-1 (CC impact in AG)
5.5 Visualization
We further visualize the ReID results with Top-5 ranking to understand how our model improves compared to the baseline [21] for aerial-to-ground person ReID in Fig. 4. Unlike the baseline which may be biased by image resolution and clothing textures, our approach pays attention to more robust features like motion patterns and body shape characteristics, which explains its successful identification of similar walking postures and body proportions despite the significant viewpoint differences between the aerial query and ground-view gallery pair. Additional examples are in Figs. 7 and 8 of the supplementary material.
6 Conclusion
We introduce AG-VPReID, a comprehensive benchmark dataset for video-based aerial-ground person ReID, addressing the critical need for a large and challenging aerial-ground dataset. We propose AG-VPReID-Net, a purpose-built three-stream person ReID framework that combines temporal-spatial processing, physics-informed normalized appearance representation, and multi-scale attention mechanisms. Our approach demonstrates state-of-the-art performance on both the AG-VPReID dataset and existing video-based ReID benchmarks. The lower performance of all approaches on the AG-VPReID dataset underscores its challenging nature and provides a robust benchmark for future research.
7 Acknowledgement
This work was supported by the Australian Research Council (ARC) Discovery Project (DP200101942) and a QUT Postgraduate Research Award. We gratefully acknowledge the Research Engineering Facility (REF) team at QUT for providing expertise and the research infrastructure essential for data collection and processing within this project.
Supplementary Material
8 Dataset
8.1 Soft-biometric Attributes
Following [31, 32], a comprehensive framework of 15 soft biometric attributes (Fig. 5) is employed to facilitate cross-view person identification. These attributes are categorized as physical traits (gender, age, height, weight, ethnicity, hairstyle, beard, and mustache) and appearance traits (glasses, head accessories, upper/lower body clothing, footwear, and accessories). This selection builds upon prior surveillance research [17, 22], prioritizing characteristics that are both discriminative and consistent across aerial and ground viewpoints for practical real-world challenges. The distribution of these soft-biometric attributes is illustrated in Fig. 6.
Figure 5: Soft-biometric attributes in our AG-VPReID dataset, showing Physical (top) and Appearance (bottom) Traits of a person from a top view. The attributes are categorized into physical characteristics (such as gender, age, height) and appearance details (such as clothing and accessories).
8.2 Long-term Re-identification
To capture realistic long-term appearance variations, we recruited 14 volunteers (consisting of nine males and five females) to participate in our data collection over a period of eight weeks with non-consecutive recording sessions. Each participant attended multiple recording sessions, with an average of four sessions per person, deliberately changing their attire between sessions. The clothing changes included variations in style (e.g., formal wear, casual wear, athletic wear), color schemes, and outer layers (e.g., jackets, coats). We instructed participants to wear clothing from their personal wardrobes to ensure naturalistic appearance variations. The recording sessions were scheduled at different times of day and under varying weather conditions, adding environmental diversity to our dataset. Each participant’s sessions were separated by a minimum interval of 14 days to maximize clothing variation and capture realistic long-term appearance changes.
Figure 6: Most common soft-biometric attributes in our dataset.
Camera $f_{x}$ $f_{y}$ $c_{x}$ $c_{y}$ $k_{1}$ $k_{2}$ $p_{1}$ $p_{2}$ Location
Bosch Outdoor 16123.85 16123.85 298.39 425.13 -0.89 1.06 0.73 1.79 (27°28’36”S, 153°01’45”E)
Bosch Indoor 17129.38 17129.38 306.20 424.81 -1.27 1.27 0.71 2.06 (27°28’40”S, 153°01’44”E)
GoPro10 17141.64 17141.64 303.62 401.73 -1.66 2.15 0.48 1.61 (27°28’35”S, 153°01’44”E)
GoPro10 16843.23 16843.23 264.75 428.89 -1.40 2.03 0.14 1.26 (27°28’35”S, 153°01’44”E)
DJI Inspire2 16467.97 16467.97 302.48 372.46 -0.93 1.14 0.41 1.68 (27°28’37”S, 153°01’46”E)
DJI M300RTK 16837.10 16837.10 291.64 418.17 -0.94 1.20 0.38 1.96 (27°28’40”S, 153°01’46”E)
Table 8: Intrinsic camera parameters and GPS coordinates
8.3 Calibration
The intrinsic camera parameters, including focal length ($f_{x}$, $f_{y}$), principal point ($c_{x}$, $c_{y}$), distortion coefficients ($k_{1}$, $k_{2}$, $p_{1}$, $p_{2}$), and GPS coordinates, are presented in Tab. 8. Relative camera locations and viewing angles are visualized in Fig. 1 of the main paper. The camera models, resolutions, lenses, and frame rates are listed in Tab. 2 of the main paper.
Stream Focus Challenges Addressed
1. Adapted Temporal-Spatial Stream Temporal shape modeling Identity-specific memory Pre-trained large vision models Motion pattern inconsistencies Temporal discontinuity Sequential feature learning
2. Normalized Appearance Stream UV map-based appearance normalization Physics-informed techniques 3D appearance representation Drastic changes in resolution Appearance variations between aerial and ground views Pose variations and partial occlusions
3. Multi-Scale Attention Stream Multi-scale feature extraction Transformer decoder Local temporal module Scale variations due to varying drone altitudes Integration of spatial and temporal information Fine-grained detail capture
Overall Framework Combination of all three streams Cross-platform visual-semantic alignment Robust person representation across different views Addresses viewpoint, resolution, scale, and occlusion
Table 9: Overview of AG-VPReID-Net streams and their contributions
9 Approach
9.1 AG-VPReID-Net Framework Overview
The AG-VPReID-Net framework, as detailed in Table 9, provides a comprehensive approach to aerial-ground video person re-identification. This framework comprises three specialized streams, each designed to address specific challenges. The Adapted Temporal-Spatial Stream focuses on motion patterns and sequential feature learning, while the Normalized Appearance Stream employs UV map-based normalization and 3D appearance representation. The Multi-Scale Attention Stream utilizes multi-scale feature extraction and a transformer decoder. These streams combine to form a robust solution. This solution handles the complex challenges of matching individuals across aerial and ground-based video footage, including viewpoint variations, occlusions, and scale differences.
9.2 Optimization
Our AG-VPReID-Net achieves optimal performance by integrating three streams. Initially, each stream produces independent feature representations. These are then combined using an adaptive weighted fusion mechanism:
$F_{combined}=\alpha F_{temporal}+\beta F_{appearance}+\gamma F_{multiscale},$ (8)
where $\alpha$, $\beta$, and $\gamma$ are learnable parameters that adapt to the input characteristics. This allows the model to dynamically emphasize different streams depending on the specific aerial-ground matching scenario.
To further improve ranking performance, we employ Reciprocal Rank Fusion (RRF) as a post-processing step. RRF combines the individual rankings from each stream to produce a final, more robust ranking:
$RRF(d)=\sum_{i=1}^{3}\frac{1}{k+r_{i}(d)},$ (9)
where $d$ is a candidate match, $k$ is a constant set to 60 in our experiments, and $r_{i}(d)$ is the rank of $d$ in the $i$-th stream. This fusion technique gives higher weight to candidates that rank highly across multiple streams, leading to improved mean Average Precision (mAP) and Rank-k accuracy in our experiments.
RRF is chosen for its robustness, handling cases where correct matches may have inconsistent rankings across streams. This is particularly important in aerial-ground ReID, where different streams may excel under different conditions (e.g., varying altitudes or occlusions). By combining rankings rather than raw scores, RRF provides a more robust final ranking that is less sensitive to individual stream failures.
10 Implementation Details
UV Map Acquisition and Processing. Our implementation pipeline uses UV maps from Texformer [41], which utilizes 3D human meshes from RSC-Net [42]. To preserve temporal relationships within individual video sequences, we integrate PhysPT [49] for more precise pose estimations. These refined poses are fed into Texformer, yielding higher-fidelity UV maps. We enhance inter-frame consistency through normalization, histogram matching, and gamma correction. The final UV map is constructed via weighted blending, combining processed UV maps with a visibility mask M = max(dot(N, V), 0), where N is the surface normal and V is the view vector. Blending weights are determined using a softmax function over mask values, ensuring smooth transitions between UV map regions.
Configuration for Stream 1: Adapted Temporal-Spatial Stream. The Adapted Temporal-Spatial Stream utilizes a pre-trained CLIP ViT-B/16 model as the visual encoder (frozen during training), complemented by a Temporal Shape Modeling (TSM) branch with 2 GRU layers ($1024$ neurons each) and an identity-aware 3D regressor. An Attention-based Shape Aggregation (ASA) module, consisting of 2 GRU layers and a self-attention mechanism, processes shape information. Temporal features are captured using a Temporal Memory Diffusion (TMD) module with multi-head self-attention. The stream is trained on 8-frame clips ($256\times 128$ resolution) with a batch size of $16$, using Adam optimizer, an initial learning rate of $5\times 10^{-3}$ with warm-up and decay, and a weight decay of $0.01$. Data augmentation includes random horizontal flipping and random erasing.
Configuration for Stream 2: Normalized Appearance Stream. The Normalized Appearance Stream processes 3D coordinates and normalized UV texture through four Omni-scale Modules. Each module incorporates a UV-space adapted Dynamic Graph Convolution (DGC) layer and three parallel branches with varying grouping rates. The network transforms the initial $m\times 6$ input (UV coordinates + RGB) into a $96\times 512$ feature map, which is then reduced to a 512-dimensional feature vector via global pooling and a fully connected layer. The training utilizes cross-entropy loss for identity classification, with Adam optimization and cosine learning rate scheduling over 1000 epochs. During inference, the final 512-dimensional feature vector serves as a robust representation of normalized appearance across multiple frames.
Configuration for Stream 3: Multi-Scale Attention Stream. The CLIP ViT-L/14 vision encoder is employed to extract multi-scale features, utilizing a Pad-and-Resize technique to uniformly adjust each frame to a resolution of $224\times 224$, thereby preserving the original body proportions. The resulting feature volume, $\mathbf{G}\in\mathbb{R}^{T\times 257\times d}$, encapsulates the embedding size of the token, $d$. Following the methods [23, 44], the feature maps from the last four layers of the image encoder (i.e., $N=4$) are utilized. Additionally, four Transformer decoder blocks ($M=4$) are applied to further process these features.
Figure 7: A2G: Comparing baseline vs our method on AG-VPReID. Green/red: correct/incorrect labels. Altitudes: 15m (orange), 120m (blue). First tracklet image is shown. Ranks show improvements in bold. Best in color.
11 Visualization
To provide a comprehensive analysis of our method’s effectiveness compared to the baseline CLIP-based approach [21], we present qualitative results through visualization examples in Fig. 7 and Fig. 8. For the aerial-to-ground (A2G) matching scenario shown in Fig. 7, we compare the retrieval results between our approach and the baseline method. Each query image is captured from aerial views at different altitudes (15m and 120m), with corresponding ground-truth matches from ground-level perspectives. The green and red boxes indicate correct and incorrect matches respectively, while improved rank positions are highlighted in bold. These visualizations demonstrate our method’s superior ability to handle severe viewpoint variations and maintain reliable person matching across aerial and ground views. Additional examples for ground-to-aerial (G2A) matching presented in Fig. 8 further validate the robustness of our approach.
Figure 8: G2A: An extended case of Fig. 7

### v2

@@
 Abstract
-We introduce AG-VPReID, a challenging large-scale benchmark dataset for aerial-ground video-based person re-identification (ReID), comprising 6,632 identities, 32,321 tracklets, and 9.6 million frames captured from drones (15-120m altitude), CCTV, and wearable cameras.
-This dataset presents a real-world benchmark to investigate the robustness of Person ReID approaches against the unique challenges of cross-platform aerial-ground settings.
-To address these challenges, we propose AG-VPReID-Net, an end-to-end framework combining three complementary streams: (1) an Adapted Temporal-Spatial Stream addressing motion pattern inconsistencies and temporal feature learning, (2) a Normalized Appearance Stream using physics-informed techniques to tackle resolution and appearance changes, and (3) a Multi-Scale Attention Stream handling scale variations across drone altitudes.
-Our approach integrates complementary visual-semantic information from all streams to generate robust, viewpoint-invariant person representations.
-Extensive experiments demonstrate that AG-VPReID-Net outperforms state-of-the-art approaches on both our new dataset and other existing video-based ReID benchmarks, showcasing its effectiveness and generalizability.
-The relatively lower performance of all state-of-the-art approaches, including our proposed approach, on our new dataset highlights its challenging nature.
-The AG-VPReID dataset, code and models are available at https://github.com/agvpreid25/AG-VPReID-Net.
+We introduce AG-VPReID, a new large-scale dataset for aerial-ground video-based person re-identification (ReID) that comprises 6,632 subjects, 32,321 tracklets and over 9.6 million frames captured by drones (altitudes ranging from 15–120m), CCTV, and wearable cameras.
+This dataset offers a real-world benchmark for evaluating the robustness to significant viewpoint changes, scale variations, and resolution differences in cross-platform aerial-ground settings.
+In addition, to address these challenges, we propose AG-VPReID-Net, an end-to-end framework composed of three complementary streams: (1) an Adapted Temporal-Spatial Stream addressing motion pattern inconsistencies and facilitating temporal feature learning, (2) a Normalized Appearance Stream leveraging physics-informed techniques to tackle resolution and appearance changes, and (3) a Multi-Scale Attention Stream handling scale variations across drone altitudes.
+We integrate visual-semantic cues from all streams to form a robust, viewpoint-invariant whole-body representation.
+Extensive experiments demonstrate that AG-VPReID-Net outperforms state-of-the-art approaches on both our new dataset and existing video-based ReID benchmarks, showcasing its effectiveness and generalizability.
+Nevertheless, the performance gap observed on AG-VPReID across all methods underscores the dataset’s challenging nature.
+The dataset, code and trained models are available at AG-VPReID-Net.
 1 Introduction
@@
 Recently, Zhang et al.
-[46] collect a synthetic dataset named CARGO, containing 108,563 images representing 5,000 identities, to complement real-world datasets.
+[46] collect a synthetic dataset named CARGO, containing 108,563 images representing 5,000 subjects, to complement real-world datasets.
 Within video-based tasks, Zhang et al.
@@
 While current aerial-ground datasets are valuable, increasing identity variation and environmental diversity would improve model robustness for real-world applications.
-In light of this, we introduce AG-VPReID, a challenging large-scale benchmark dataset for Aerial-Ground Video-based Person ReID.
-AG-VPReID comprises 6,632 identities, 32,321 tracklets, and approximately 9.6 million frames, captured across multiple dates and times of day using a combination of three platforms: aerial drones operating at various altitudes (15-120m), stationary CCTV cameras and wearable mobile cameras.
+In light of this, we introduce AG-VPReID, a comprehensive large-scale benchmark dataset for Aerial-Ground Video-based Person ReID.
+AG-VPReID comprises 6,632 subjects, 32,321 tracklets, and over 9.6 million frames, captured across multiple dates and times of day using a combination of three platforms: aerial drones operating at various altitudes (15-120m), stationary CCTV cameras and wearable mobile cameras.
 This dataset significantly surpasses existing video-based ReID benchmarks in terms of scale, diversity, and real-world applicability with the highest number of identities, the highest number of tracklets, the highest drone flying altitudes, and the most diverse platforms.
 The key characteristics of AG-VPReID include: drastic view changes between aerial and ground perspectives; a large number of annotated identities across multiple sessions; rich outdoor scenarios with varying environmental conditions; significant differences in resolution between aerial and ground footage; and both controlled scenarios with clothing changes and in-the-wild pedestrian traffic.
-Aerial-ground person re-identification presents unique challenges due to significant appearance variations between aerial and ground-level views.
+Aerial-ground person ReID presents unique challenges due to significant appearance variations between aerial and ground-level views.
 These variations include extreme viewpoint differences, drastic changes in resolution and scale, partial occlusions, and temporal discontinuities caused by high-flying altitudes and long-range captures.
 Traditional video-based person ReID methods, although effective in ground-based settings [26, 45], often struggle in aerial-ground scenarios due to the complex combination of inconsistent motion patterns and the aforementioned variations.
-To address these limitations, we introduce AG-VPReID-Net, an end-to-end framework for Aerial-Ground Video-based Person Re-Identification.
+To address these challenges, we introduce AG-VPReID-Net, an end-to-end framework for Aerial-Ground Video-based Person Re-Identification.
 Unlike existing state-of-the-art methods focused on single-view or ground scenarios, AG-VPReID-Net features three complementary streams tailored for aerial-ground challenges: i) An Adapted Temporal-Spatial Stream enhances traditional temporal modeling by integrating identity-specific memory and temporal shape analysis.
@@
 (3) AG-VPReID-Net sets new state-of-the-art performance on the AG-VPReID and existing video-based ReID benchmarks, demonstrating our approach’s effectiveness and generalizability across different settings.
-Figure 1: Our dataset was captured using six diverse cameras (aerial drones, CCTVs, GoPros) with sample images and camera locations shown on the right.
-Left shows cross-camera appearance variations of two pedestrians across different sessions and times of the day.
+Figure 1: Our AG-VPReID dataset was captured using a variety of six cameras, including aerial drones, CCTVs, and GoPros.
+Sample images and camera locations are illustrated on the right side of the figure.
+The left side depicts the cross-camera appearance variations of two pedestrians, showcasing differences across various sessions and times of the day.
 2 Prior Work
@@
 1 compares our AG-VPReID dataset with others across multiple dimensions.
-Video-based Person ReID Approaches.
+Video-based Person ReID.
 Video-based person ReID methods have evolved to leverage both spatial and temporal cues.
@@
 3 AG-VPReID Dataset
-This section provides an overview of our AG-VPReID dataset creation process.
-We describe our methods for collecting video footage in Sec.
+This section offers a detailed overview of the creation process for our AG-VPReID dataset.
+We describe the methods used for collecting video footage in Sec.
 3.1.
 Sec.
-3.2 outlines our annotation procedures.
+3.2 introduces our annotation procedures.
 Sec.
-3.3 compares AG-VPReID with existing datasets, emphasizing its unique features.
+3.3 compares AG-VPReID with existing datasets, highlighting its unique features.
 3.1 Dataset Collection
-The AG-VPReID dataset was collected over 20 days, comprising 10 morning sessions (8:30am-10:00am) and 10 afternoon sessions (3:30pm-5:00pm), each lasting 60 minutes.
-We captured data using two drones, two CCTV cameras, and two wearable cameras.
-Drones flew at four altitudes (15m, 30m, 80m, 120m) for 15 minutes each per session.
-The dataset contains 240 hours of video footage in total.
-Tab.
-2 details the equipment specifications.
-The dataset includes varying resolutions, frame rates, and perspectives, ranging from ground level to 120-meter aerial views.
+The AG-VPReID dataset was collected over a period of 20 days, including 10 morning sessions (8:30am-10:00am) and 10 afternoon sessions (3:30pm-5:00pm), with each session lasting 60 minutes.
+Data capture involved two drones, two CCTV cameras, and two wearable cameras.
+Each drone operated at four different altitudes—15m, 30m, 80m, and 120m—for 15 minutes per session, providing a comprehensive range of aerial views.
+In total, the dataset comprises 240 hours of video footage, documented in Tab.
+2, which includes detailed specifications of the equipment used.
+The dataset features diverse resolutions, frame rates, and perspectives, extending from ground level to 120-meter aerial views.
 Fig.
@@
 Table 2: Equipment specifications for the AG-VPReID dataset.
-To ensure professional drone operations, we hired a specialized team consisting of one Remotely Piloted Aircraft Systems (RPAS) engineer, one Chief Remote Pilot, and one RPAS technician.
-This team was present for all 20 data collection days, responsible for setting up and flying the drones, capturing high-quality aerial footage, and performing initial data preprocessing.
-Their expertise was crucial in maintaining consistency and quality across the dataset, particularly for the challenging aerial perspectives.
+To ensure professional drone operations, a specialized team (one RPAS engineer, one Chief Remote Pilot, one RPAS technician) managed all 20 data collection days, handling flights, capturing aerial footage, and performing initial data preprocessing.
 3.2 Labeling Process
 The AG-VPReID dataset uses YOLOv8x for person detection and tracking [16], extracting images from all frames across multiple cameras.
-It includes both short-term and long-term ReID scenarios, with participants changing clothing for long-term testing.
-Identity matching was performed by multiple expert annotators with research assistant help to ensure accuracy and consistency.
-Following [32], we manually labeled 15 selective soft-biometrics attributes per identity to enhance attribute-based person ReID capabilities.
-For the attribute list, see 8.1.
+It includes both short-term and long-term ReID scenarios, the latter including instances where participants change clothes to test long-term identity persistence.
+Identity matching was performed by a team of expert annotators, supported by research assistants, to ensure both accuracy and consistency in the matching process.
+Following [32], we manually annotated each identity with 15 selective soft-biometrics attributes to enhance the dataset’s utility for attribute-based person ReID applications.
+For a detailed list of these attributes, refer to 8.1.
 3.3 Dataset Characteristics
@@
 1) The highest number of identities and tracklets.
-With 6,632 unique identities and 32,321 tracklets, our dataset significantly exceeds existing ones.
-It has almost double the identities of ground-based datasets like LSVID [18] (3,772) and over six times the tracklets of G2A-VReID [47] (5,576).
-With 9.6 million frames, it provides 50 times more data than other aerial-ground datasets, though MEVID [3] has slightly more frames but fewer identities and tracklets.
+AG-VPReID sets a new benchmark with 6,632 unique identities and 32,321 tracklets, substantially surpassing existing datasets.
+It holds nearly twice as many identities as prominent ground-based datasets such as LSVID [18], which contains 3,772 identities, and more than six times the tracklets of the next largest aerial-ground dataset, G2A-VReID [47], which includes 5,576 tracklets.
+Additionally, AG-VPReID encompasses over 9.6 million frames, providing 50 times the volume of frames compared to other aerial-ground datasets.
+While MEVID [3] exceeds in frame count with 10.46 million, it does not match AG-VPReID in terms of identity and tracklet numbers.
 2) The most diverse platforms.
 Our AG-VPReID dataset is the first to incorporate aerial, ground, and wearable platforms for video-based person ReID.
-The inclusion of wearable cameras introduces a novel dimension with high-quality first-person perspectives.
+The inclusion of wearable cameras provides a novel dimension with high-quality first-person perspectives.
 This combination results in extreme variations in resolution and subject size across platforms: UAV (18$\times$37 to 293$\times$542 pixels), CCTV (22$\times$23 to 172$\times$413 pixels), and GoPro (25$\times$48 to 483$\times$1078 pixels).
 3) The highest flying altitudes.
-AG-VPReID features footage from altitudes up to 120 meters, exceeding previous datasets’ 60-meter maximum [47].
+AG-VPReID features footage from altitudes reaching up to 120 meters, exceeding existing datasets’ 60-meter maximum [47].
 This introduces challenges: 1) extreme viewpoints with perspective distortions; 2) multiple scales with varying resolutions; and 3) image quality issues (Fig.
@@
 8 for details.
-Figure 2: Main challenges in AG-VPReID: extreme viewpoints, varying resolutions and sizes, pose/illumination variations, occlusions, and similar clothing.
+Figure 2: The AG-VPReID dataset presents several key challenges: extreme viewpoints, varying resolutions and subject sizes, pose/illumination variations, occlusions, and similar clothing among subjects.
 3.4 Ethics and Privacy
-The research has been ethics-approved by related authorities for data collection and usage of the collected data for research purposes.
-We use “Deface” [5] to blur facial regions in media to protect individual identities.
-Data is securely stored with authorized access only.
-For volunteers involved in the study, ethical clearance was obtained, and informed consent was secured prior to their participation.
-Details on privacy protection and data usage are available at our project repository.
-4 Approach
+This research received ethics approval for data collection and usage.
+We implement “Deface” [5] to blur faces, secure data storage, and obtained informed consent from all participants.
+Details are available at our project repository.
+4 AG-VPReID-Net
 We propose AG-VPReID-Net, a purpose-built framework addressing aerial-ground ReID’s unique challenges.
@@
 By combining a lightweight Transformer decoder with a local temporal module, this approach dynamically integrates spatial and temporal information, thereby enhancing our framework’s ability to accurately capture essential person-specific details.
-Specifically, for each frame $\mathcal{I}_{t}$ within the sequence $\mathcal{V}{y_{i}}$, the CLIP vision encoder CLIP [36] is employed to extract features independently.
+Specifically, for each frame $\mathcal{I}_{t}$ within the sequence $\mathcal{V}{y_{i}}$, the CLIP vision encoder [36] is employed to extract features independently.
 The process collects tokens from various layers at regular intervals to compile a detailed feature map that captures spatial correspondences.
@@
 5.1 Datasets and Evaluation Metrics
-We evaluate our method on AG-VPReID and four established video-based person ReID datasets: iLIDS [38], Mars [51], LS-VID [18] and G2A-VReID [47].
-For AG-VPReID, 3,013 identities with ground and aerial views were split 1:1 for training and testing [31, 32].
-Training and testing details are presented in Tab.
-3.
-We use Cumulative Matching Characteristic (CMC) at Rank-1 and mean Average Precision (mAP) for evaluation.
+We conducted evaluations of our method using the AG-VPReID and four established video-based person ReID datasets: iLIDS [38], Mars [51], LS-VID [18] and G2A-VReID [47].
+For AG-VPReID, we used a balanced split of 3,013 identities with both ground and aerial views, dividing them equally for training and testing purposes [31, 32].
+Details on the training and testing configurations are provided in Table 3.
+We evaluate performance using the Cumulative Matching Characteristic (CMC) at Rank-1 and the mean Average Precision (mAP).
 Case Subset # IDs # Tracklets # Images (M)
@@
 The architecture consists of three streams: an Adapted Temporal-Spatial Stream (CLIP ViT-B/16), a Normalized Appearance Stream for 3D coordinates and UV textures, and a Multi-Scale Attention Stream (CLIP ViT-L/14).
-Implementation details are in the supplementary material Sec.
-10.
+More implementation details can be found in the supplementary.
 5.3 Comparison with State-of-the-Art Methods
@@
 Ground-to-Ground Datasets.
-Our method achieves superior performance on MARS (91.5% mAP, 93.2% Rank-1), exceeding CLIP-ReID by 3.4% mAP.
+Our method achieves superior performance on MARS (91.5% mAP, 93.2% Rank-1), outperforming CLIP-ReID by 3.4% mAP.
 On LS-VID (87.3% mAP, 93.2% Rank-1), we surpass LSTRL by 4.9% mAP.
-For iLIDS-VID, we reach 96.3% Rank-1, outperforming MFA by 3.0%.
+For iLIDS-VID, we reach 96.3% Rank-1, which is 3.0% higher than MFA.
 Cross-Platform Datasets.
@@
 Tab.
-5 shows St-1 achieves the strongest individual performance (71.52% A2G, 74.80% G2A Rank-1).
+5 shows that St-1 achieves the strongest individual performance (71.52% A2G, 74.80% G2A Rank-1).
 St-2 and St-3 show moderate results (58.40% and 61.65% A2G Rank-1).
-Combined streams demonstrate complementary strengths, with St-123 achieving the best results (71.91% A2G, 75.57% G2A Rank-1) by integrating all three streams.
+Combined streams demonstrate complementary strengths, with St-123 achieving the best results (71.91% A2G, 75.57% G2A Rank-1) by integrating the three streams.
 Figure 4: Baseline vs our method on AG-VPReID dataset. Green/red: correct/incorrect labels. First tracklet image shown. Ranks show improvements in bold.
 Impact of Altitude.
-Table 6 shows performance decreases with altitude, dropping most between 30m and 80m.
-A2G Rank-1 drops  11% across streams due to scale variations and reduced resolution.
-St-1 shows remarkable robustness at high altitudes versus other streams (52.47% vs 38.12%/35.13% at 120m A2G), demonstrating a +17.34% improvement through our temporal modeling approach.
-Similarly, St-2’s physics-informed UV mapping provides a +7.57% Rank-1 improvement at 120m (42.70% vs.
-35.13%), while St-3’s multi-scale attention design yields a +17.72% improvement (52.85% vs.
-35.13%) in the same setting.
-The poor high-altitude performance by all methods highlights unique challenges not present in typical low-altitude datasets.
-The St-123 maintains best performance by leveraging complementary stream strengths across all altitudes.
+Table 6 shows performance decreasing with altitude, most significantly between 30m-80m.
+A2G Rank-1 drops $\sim$11across streams.
+At 120m, St-1 demonstrates robustness (52.47% vs 38.12%/35.13%), achieving +17.34% improvement through temporal modeling.
+St-2’s physics-informed UV mapping provides +7.57% improvement (42.70% vs.
+35.13%), while St-3’s multi-scale attention yields +17.72% improvement (52.85% vs.
+35.13%).
+St-123 maintains best performance across all altitudes by combining stream strengths.
 Clothing Changes vs.
@@
 See Table 7.
-Table 7: Impact of clothing changes (CC) vs. camera angles.
+5.5 Visualization
+We further visualize the ReID results with Top-5 ranking to understand how our model improves compared to the baseline [21] for aerial-to-ground person ReID in Fig.
+4.
+Unlike the baseline which may be biased by image resolution and clothing textures, our approach pays attention to more robust features like motion patterns and body shape characteristics, which explains its successful identification of similar walking postures and body proportions despite the significant viewpoint differences between the aerial query and ground-view gallery pair.
+Additional examples are in Figs.
+7 and 8 of the supplementary material.
 Scenario Rank-1 mAP Key Observation
@@
 AG-DiffClothes 65.83 57.52 $-6.08\%$ Rank-1 (CC impact in AG)
-5.5 Visualization
-We further visualize the ReID results with Top-5 ranking to understand how our model improves compared to the baseline [21] for aerial-to-ground person ReID in Fig.
-4.
-Unlike the baseline which may be biased by image resolution and clothing textures, our approach pays attention to more robust features like motion patterns and body shape characteristics, which explains its successful identification of similar walking postures and body proportions despite the significant viewpoint differences between the aerial query and ground-view gallery pair.
-Additional examples are in Figs.
-7 and 8 of the supplementary material.
+Table 7: Impact of clothing changes (CC) vs. camera angles.
 6 Conclusion
-We introduce AG-VPReID, a comprehensive benchmark dataset for video-based aerial-ground person ReID, addressing the critical need for a large and challenging aerial-ground dataset.
-We propose AG-VPReID-Net, a purpose-built three-stream person ReID framework that combines temporal-spatial processing, physics-informed normalized appearance representation, and multi-scale attention mechanisms.
-Our approach demonstrates state-of-the-art performance on both the AG-VPReID dataset and existing video-based ReID benchmarks.
-The lower performance of all approaches on the AG-VPReID dataset underscores its challenging nature and provides a robust benchmark for future research.
+We introduce AG-VPReID, a comprehensive dataset for video-based aerial-ground person ReID, addressing the critical need for a large and challenging aerial-ground dataset.
+We also propose AG-VPReID-Net, a purpose-built three-stream person ReID framework that combines temporal-spatial processing, physics-informed normalized appearance representation, and multi-scale attention mechanisms.
+This approach achieves state-of-the-art performance on both the AG-VPReID dataset and existing video-based ReID benchmarks.
+Notably, the relatively lower performance across all approaches on AG-VPReID highlights its demanding nature and establishes it as a robust benchmark for advancing future research in the field.
 7 Acknowledgement